OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10

OpenStudy (anonymous):

!)x + 4y + z = –10 x + 2y + 2z = –10 => 2)-x -2y-2z = 10 from 1 and 2 we conclude that : 2y - z = 0 ** then we have : x + 4y + z = –10=>1) -3x-12y-3z = 30 2) 3x – 3y + 6z = –21 from 1 and 2 again we conclude that :-15y +3z = 9 or -5y + z = 3 *** then by solving *** and ** we conclude that y = 1 and so on ...

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

any question I can help you with?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = –7 x – 3y + 4z = –14 x – 2y – 3z = –11

OpenStudy (anonymous):

In the following equation System I use the deletion method(I 'm not sure about the name because of my poor English)for the second and third equations.this way the X can be omitted easily and I will have an equation with y and z only : x-3y+4z = -14 => -x +3y-4z = 14 (1) x – 2y – 3z = –11 (2) then from 1 and 2 we conclude that : y - 7z = 3 ** we do the same procedure for the either first and second equation or first and third equation. let's choose first and second equation for example : in the second equation we have x – 3y + 4z = –14 and you see the coefficient of x is 2 so we should multiply this equation to (-2) because we want to omit x and find a new equation between y and z : we have : 2x + y + z = –7 (1) x – 3y + 4z = –14 => -2x +6y -8z = 28 (2) from 1 and 2 we conclude that : 7y - 7z = 21 *** then we can solve the *** and ** equations to find z and y.

OpenStudy (anonymous):

let me know if you get what I said or not, dear...